OpenStudy (kaioken360):

f(x)=4x^7 and g(x)=6x^2. What is f(1)/g(2)? A) 1 B) 1/6 C) 1/4 D) 1/24

OpenStudy (anonymous):

Are there any x's in the 4^7 and 6^2?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

So for this, you are given the values you will input to solve.

OpenStudy (anonymous):

You input 1 for the x in the f(x) problem and input 2 for the x in the g(x) problem.

OpenStudy (anonymous):

After that you place the answers over each other with the numerator being the f(x) problem.

OpenStudy (anonymous):

Do you want to attempt it kai?

Directrix (directrix):

To find f(1), replace x in f(x) = 4x^7 by 1 f(1) = 4 (1)^7 = ? @kaioken360 Post what you get.

OpenStudy (anonymous):

\[f(x)=4x^7~and~g(x)=6x^2\]substitute the value into\[\frac{ f(1) }{ g(2) }\]\[\frac{ f(1) }{ g(2) }=\frac{ 4(1)^7 }{ 6(2)^2 }\]\[\frac{ f(1) }{ g(2) }=\frac{ 4\times1 }{ 6\times2\times2 }\]\[\frac{ f(1) }{ g(2) }=\frac{ 4 }{ 24 }\]Then,try to reduce the result. We noticed that 4 and 24 can be divide by 4.\[\frac{ f(1) }{ g(2) }=\frac{ 1\times4 }{ 6\times4 }\]\[\frac{ f(1) }{ g(2) }=\frac{ 1\times\cancel{4} }{ 6\times\cancel{4} }\]\[Therefore,\frac{ f(1) }{ g(2) }=?\]solve it @kaioken360